OpenStudy (anonymous):

Solve (x + 2)(2x + 3) = 6

OpenStudy (anonymous):

(x+2)(2x+3)=6 =x*(2x+3)+2*(2x+3)=6 =(2x^2+3x)+(4x+6)=6 =2x^2+7x+6=6 =2x^2=-7x

OpenStudy (anonymous):

looks right :) just to finish it off, that means 2x=-7, so x will equal -3.5. However, because x does need to have two roots, the second root will be zero, because in the step that Syamantak left off in, we have that 2x^2=-7x, so 0 is a solution to that equation as well.

OpenStudy (anonymous):

I have 3 answers which one is the correct one {-2,-3/2}, {-7/2, 0}, {0}

OpenStudy (anonymous):

the second one is the correct choice, (-7/2,0)